OpenStudy (anonymous):

\[\log_{5}\log_{3}\log_{2}x = 0 \] Find x

OpenStudy (anonymous):

\[\log _3\log _2x=5^0\]

OpenStudy (anonymous):

\[\log_{5} (\log_{3}\log_{2}x) =0 \] \[\log_{3}\log_{2}x = 1 \]

OpenStudy (anonymous):

\[\log_2x=3\]

OpenStudy (anonymous):

\[\log_{2}x = 3 \] \[x=8\]

OpenStudy (anonymous):

@Yahoo! you are just reading mind

OpenStudy (anonymous):

lol.....Thxx

OpenStudy (anonymous):

easy one this time...